OpenStudy (anonymous):

Solve –3x2 + 30x – 90 = 0

OpenStudy (mertsj):

Factor out -3 first.

OpenStudy (anonymous):

x = -30+ 30i / -6 = 5 - 30i and x= 5+30i

OpenStudy (mertsj):

\[-3(x^2-10x+30)=0\]

OpenStudy (anonymous):

the roots will be complex...it can't be factorized..

OpenStudy (mertsj):

\[x=\frac{10\pm \sqrt{100-4(30)}}{2}=\frac{10\pm \sqrt{100-120}}{2}\]

OpenStudy (mertsj):

\[x=\frac{10\pm \sqrt{-20}}{2}=\frac{10\pm 2\sqrt{5}i}{2}=5\pm \sqrt{5}i\]

OpenStudy (mertsj):

You forgot to divide the 10 by 2 lilmagda105

OpenStudy (anonymous):

no i gave you the options of answers

OpenStudy (mertsj):

There is only one answer on that post and it is incorrect.

OpenStudy (mertsj):

The answers I gave you are correct.

OpenStudy (anonymous):

 is that the answer